# Towards Scalable Language-Image Pre-training for 3D Medical Imaging

### v2

@@ 5 Conclusion
  • D: Complexity
  • E: Discussion
@@ Appendix A Dataset
As mentioned in Section 3.2, this section provides additional details on the two datasets we collected: BrainMRI220K and HeadCT240K. For other datasets, such as CT-RATE [3], Rad-ChestCT [24], BraTS 2023 [59, 60, 47, 48, 49, 50], NYU-Mets [52], UCSF-Mets [51], RSNA [22], and CQ500 [21], we refer readers to their original publications. Additionally, we include a discussion of our preprocessing strategy.
@@ A.3 Discussion
To the best of our knowledge, this is the first work to handle 3D medical datasets at the scale of BrainMRI220K and HeadCT240K, which comprise millions of scans. Our preprocessing strategy, introduced in Section 3.2, follows standard practices where possible, but also diverges due to the scale and diversity of our dataset. Specifically, percentile clipping for brain MRI and HU value truncation for head CT follow standard practices [61, 20]. For head CT, we do not concatenate scans with different HU truncations along the channel dimension, as our method is capable of processing multiple scans simultaneously. However, our strategy regarding orientation and spacing may differ from established preprocessing pipelines [62, 4, 3, 6]. Here, we provide a discussion to justify the rationale behind our approach.
@@
Orientation. CT and MRI scans are acquired using different planes, including axial, coronal, and sagittal. Isotropic data are rare in routine radiology studies; the through-plane spacing is typically larger than the in-plane spacing. In the context of batch construction, where all samples in a batch must share the same shape, standardizing the orientation can lead to downsampling along the in-plane axes and upsampling along the through-plane axis. Upsampling in pixel space does not introduce new information, whereas downsampling results in information loss. From the perspective of data augmentation, the orientation can change when 3D random rotation is applied [62]. So, why do we standardize the orientation at the beginning then augment it during training? Instead, we consistently align the first dimension (depth) with the through-plane axis. Please note that the in-plane orientation is still standardized as RP, RI, or PI. We treat the diversity of acquisition planes (as shown in Figure 7(c) and Figure 9(c)) in our dataset as a form of natural data augmentation.
@@
Spacing. Standardizing spacing across the entire dataset is a common practice in medical image segmentation to ensure that convolution filters interpret anatomical structures consistently [62]. However, HyperSpace [63] demonstrates that, with appropriate spacing augmentation, convolutional networks can learn spacing variations. Therefore, architectures like Vision Transformers, which can learn scale-invariant features [11], should also be capable of learning spacing-invariant features with sufficient spacing augmentation. This is because spacing invariance is a direct extension of scale equivariance from 2D to 3D space. Given the diversity of spacing settings (as shown in Figure 7(b) and Figure 9(b)) in our dataset, we believe it naturally occupies sufficient spacing augmentation for the model to learn spacing-equivariant features. Therefore, when constructing a batch, we first resize each volume to (48,256,256), and then apply a center crop to obtain (48,224,224). Here, 48 represents the median number of slices, which does not substantially distort the original scan.
@@
$M$ scans per study. As the number of scans per study varies, we randomly sample $M=10$ scans per study at each training step to construct a batch. An equal number of positional embeddings is also sampled for these $M$ scans from a total of $M_{\text{max}}=40$ positional embeddings at each step. This means the model is exposed to $M_{\text{max}}$ different positions over the course of training, while observing $M$ positions at each step. This is analogous to a dropout strategy, with a scan dropout rate of $1-\frac{M}{M_{\text{max}}}$. During evaluation, the model is able to process studies containing up to $M_{\text{max}}$ scans without the need for interpolation.
@@ Appendix B Implementation
In this section, we summarize the model configuration described in Section 3.2 and provide details of the pre-training setup referenced in both Section 3.2 and Section 4.3. For linear probing evaluation on head CT, we refer readers to FM-HeadCT [20].
@@ B.1 Chest CT
Table 3: Model (Chest CT).
@@
Text Encoder CXR-BERT [7]
context length 512 [3]
Table 4: Pre-training (Chest CT).
@@
truncate (window) [-1150,350]
normalize (mean, std) [0.449,0.226]
@@
optimizer AdamW [64]
base learning rate 1e-5
@@
optimizer momentum $\beta_{1}$, $\beta_{2}$ = 0.9, 0.98 [2]
learning rate schedule cosine decay [65]
warmup (in steps) 47 [6]
@@
Table 5: Evaluation (Chest CT).
CT-RATE [3] Rad-ChestCT [24]
@@
The model architecture and pre-training configuration for chest CT are summarized in Table 4 and Table 4, respectively. These configurations closely follows prior works such as CT-CLIP [3], OpenCLIP [45], and BiomedCLIP [32]. In the pre-training configuration, we report the base learning rate corresponding to a batch size of 64. During training, we follow the linear learning rate scaling rule [19]: $lr=base\_lr\times\frac{batch\_size}{64}$.
The diseases used for zero-shot evaluation are listed in Table 5. This setup is consistent with fVLM [6] and all methods reported in Table 2.
@@ B.2 Brain MRI/Head CT
Table 6: Model (Brain MRI/Head CT).
@@
Text Encoder PubMedBERT [44, 32]
context length 256 [32]
Table 7: Pre-training (Brain MRI/Head CT).
@@
truncate (percentile) [0.05,0.95]
normalize (mean, std) [0.449,0.226]
report GPT-3.5
optimizer AdamW [64]
base learning rate 1e-4
@@
optimizer momentum $\beta_{1}$, $\beta_{2}$ = 0.9, 0.95 [19]
learning rate schedule cosine decay [65]
warmup (in steps) 2000
@@
The model architecture and pre-training configuration for brain MRI/head CT are summarized in Table 7 and Table 7, respectively. These configurations also follows prior works [32, 45, 19]. In the pre-training configuration, we report the base learning rate corresponding to a batch size of 128. During training, we follow the linear learning rate scaling rule [19]: $lr=base\_lr\times\frac{batch\_size}{128}$.
@@ Appendix C Prospective Evaluation
Appendix D Complexity
D.1 Analysis
@@
D.2 Experiments
We conduct experiments comparing ViT [11], the 3D Swin Transformer implemented in MONAI [66], and our HLIP visual encoder. Since Swin is not compatible with multi-scan inputs, the experiment is conducted on a single-channel 3D input of shape (B,1,224,224,224), where B=1 denotes the batch size. Both ViT and HLIP use a ViT-Base architecture with patch_size=(16,16,16). The Swin Transformer is configured with dim=128, patch_size=(4,4,4), window_size=(7,7,7), depth=(2,2,12,2), and head=(4,8,16,32). Following the original Swin Transformer [12] while matching the model size.
@@ D.2 Experiments
model size
throughput
(img/s)
$\uparrow$=better
memory
$\downarrow$=better
@@ Appendix E Discussion
Moreover, we observe that zero-shot transferability does not always correlate with the number of studies used for training. For example, although keyword search yields more patients with meningioma or metastasis than with glioma, zero-shot performance on glioma is substantially higher. While tumor size may partially explain this discrepancy, we believe it warrants further investigation in future work.

### v3

@@ 5 Conclusion
  • D: Statistical Analysis
  • E: Complexity
  • F: Discussion
@@ Appendix A Dataset
As mentioned in Section 3.2, this section provides additional details on the two datasets we collected: BrainMRI220K and HeadCT240K. For other datasets, such as CT-RATE (Hamamci et al., 2024), Rad-ChestCT (Draelos et al., 2021), BraTS 2023 (Menze et al., 2014; Bakas et al., 2017; Baid et al., 2021; LaBella et al., 2023; Moawad et al., 2024; Kazerooni et al., 2024), NYU-Mets (Link et al., 2024), UCSF-Mets (Rudie et al., 2024), RSNA (Flanders et al., 2020), and CQ500 (Chilamkurthy et al., 2018), we refer readers to their original publications. Additionally, we include a discussion of our preprocessing strategy.
@@ A.3 Discussion
To the best of our knowledge, this is the first work to handle 3D medical datasets at the scale of BrainMRI220K and HeadCT240K, which comprise millions of scans. Our preprocessing strategy, introduced in Section 3.2, follows standard practices where possible, but also diverges due to the scale and diversity of our dataset. Specifically, percentile clipping for brain MRI and HU value truncation for head CT follow standard practices (Ma et al., 2024; Zhu et al., 2025). For head CT, we do not concatenate scans with different HU truncations along the channel dimension, as our method is capable of processing multiple scans simultaneously. However, our strategy regarding orientation and spacing may differ from established preprocessing pipelines (Isensee et al., 2021; Blankemeier et al., 2024; Hamamci et al., 2024; Shui et al., 2025).
@@
Orientation. CT and MRI scans are acquired using different planes, including axial, coronal, and sagittal. Isotropic data are rare in routine radiology studies; the through-plane spacing is typically larger than the in-plane spacing. In the context of batch construction, where all samples in a batch must share the same shape, standardizing the orientation can lead to downsampling along the in-plane axes and upsampling along the through-plane axis. Upsampling in pixel space does not introduce new information, whereas downsampling results in information loss. From the perspective of data augmentation, the orientation can change when 3D random rotation is applied (Isensee et al., 2021). So, why do we standardize the orientation at the beginning then augment it during training? Instead, we consistently align the first dimension (depth) with the through-plane axis. Please note that the in-plane orientation is still standardized as RP, RI, or PI. We treat the diversity of acquisition planes (as shown in Figure 7(c) and Figure 9(c)) in our dataset as a form of natural data augmentation.
@@
Spacing. Standardizing spacing across the entire dataset is a common practice in medical image segmentation to ensure that convolution filters interpret anatomical structures consistently (Isensee et al., 2021). However, HyperSpace (Joutard et al., 2024) demonstrates that, with appropriate spacing augmentation, model can learn spacing-invariant features with sufficient spacing augmentation. Given the diversity of spacing settings (as shown in Figure 7(b) and Figure 9(b)) in our dataset, we believe it naturally occupies sufficient spacing augmentation for the model to learn spacing-invariant features. Therefore, when constructing a batch, we first resize each volume to (48,256,256), and then apply a center crop to obtain (48,224,224). Here, 48 represents the median number of slices, which does not substantially distort the original scan.
@@
M scans per study. As the number of scans per study varies, we randomly sample $M\texttt{=}10$ scans per study at each training step to construct a batch. An equal number of positional embeddings is also sampled for these $M$ scans from a total of $M_{\text{max}}\texttt{=}40$ positional embeddings at each step. This means the model is exposed to $M_{\text{max}}$ different positions over the course of training, while observing $M$ positions at each step. This is analogous to a dropout strategy, with a scan dropout rate of $1-\frac{M}{M_{\text{max}}}$. During evaluation, the model is able to process studies containing up to $M_{\text{max}}$ scans without the need for interpolation.
@@ Appendix B Implementation
In this section, we summarize the model configuration described in Section 3.2 and provide details of the pre-training setup referenced in both Section 3.2 and Section 4.3. For linear probing evaluation on head CT, we refer readers to FM-HeadCT (Zhu et al., 2025).
@@ B.1 Chest CT
Table 3: Model (Chest CT)
@@
Text Encoder CXR-BERT (Boecking et al., 2022)
context length 512 (Hamamci et al., 2024)
Table 4: Pre-training (Chest CT)
@@
window [-1150,350]
mean, std [0.449,0.226]
@@
optimizer AdamW (Loshchilov and Hutter, 2017)
$\beta_{1}$, $\beta_{2}$ 0.9, 0.98 (Radford et al., 2021)
base lr 1e-5
@@
lr schedule cosine (Loshchilov and Hutter, 2016)
warmup steps 47 (Shui et al., 2025)
@@
Table 5: Evaluation (Chest CT)
CT-RATE (Hamamci et al., 2024) Rad-ChestCT (Draelos et al., 2021)
@@
The model architecture and pre-training configuration for chest CT are summarized in Table 4 and Table 4, respectively. These configurations closely follows prior works such as CT-CLIP (Hamamci et al., 2024), OpenCLIP (Cherti et al., 2023), and BiomedCLIP (Zhang et al., 2023a). In the pre-training configuration, we report the base learning rate corresponding to a batch size of 64. During training, we follow the linear learning rate scaling rule (Li et al., 2023): $lr=base\texttt{\_}lr\times\frac{batch\texttt{\_}size}{64}$.
The diseases used for zero-shot evaluation are listed in Table 5. This setup is consistent with fVLM (Shui et al., 2025) and all methods reported in Table 2.
@@ B.2 Brain MRI/Head CT
Table 6: Model (Brain MRI/Head CT)
@@
Text Encoder PubMedBERT (Gu et al., 2020)
context length 256 (Zhang et al., 2023a)
Table 7: Pre-training (Brain MRI/Head CT)
@@
truncate [0.05,0.95]
mean, std [0.449,0.226]
report Prima (Lyu et al., 2026)
optimizer AdamW (Loshchilov and Hutter, 2017)
$\beta_{1}$, $\beta_{2}$ 0.9, 0.95 (Li et al., 2023)
base lr 1e-4
@@
lr schedule cosine (Loshchilov and Hutter, 2016)
warmup steps 2000
@@
The model architecture and pre-training configuration for brain MRI/head CT are summarized in Table 7 and Table 7, respectively. We process the reports using the method proposed in Prima (Lyu et al., 2026), which leverages a large language model to summarize the original radiologist reports and reduce biases related to writing style, grammar, and wording habits. Given the rapid development of modern language models, we do not consider this fully automatic text processing applied to existing radiologist reports to be equivalent to the data curation pipeline on the visual side, which relies on radiologists and imposes additional burdens.
@@ Appendix C Prospective Evaluation
Appendix D Statistical Analysis
We investigate statistical metrics for HLIP and the original ViT on the Pub-Brain-5 dataset. First, using the zero-shot prompt reported in the paper, "This brain MRI shows: {disease}.", we find that the p-value between the two models is less than 0.0001. We then perform inference using four additional prompts: "This brain MRI shows: likely {disease}.", "This brain MRI shows: {disease} identified.", "This brain MRI shows: suggesting {disease}.", and "This brain MRI shows: compatible with {disease}." Under the five prompts considered, the 5-way disease classification performance is $63.4_{\pm 3.8}$ for HLIP and $40.6_{\pm 4.0}$ for ViT. This result demonstrates a clear benefit of the HLIP methodology over the original ViT.
Table 8: Results of zero-shot classification on Pub-Brain-5. We report the balanced accuracy of HLIP under five prompts.
Anomaly Detection Tumor Disease
Stroke Glioma Meningioma Metastasis mean
Pub-Brain-5 $86.9\pm 6.0$ $92.4\pm 4.8$ $76.1\pm 3.4$ $84.9\pm 4.9$ $85.4\pm 3.7$ $62.3\pm 1.4$ $63.4\pm 3.8$
Pub-Brain-5-GT $90.4\pm 5.2$ $92.5\pm 5.0$ $6.3\pm 3.3 77$$7.4\pm 6.9$ $84.1\pm 3.7$ $53.8\pm 0.9$ $59.3\pm 3.7$
In Table 8, we additionally report the full results of HLIP under this setting on the Pub-Brain-5 and Pub-Brain-5-GT benchmark, for reference in future work.
Appendix E Complexity
In this section, we analyze the computational efficiency of the proposed hierarchical attention mechanism from both theoretical and practical perspectives.
E.1 Analysis
@@
E.2 Experiments
We conduct experiments comparing ViT (Dosovitskiy et al., 2020), the 3D Swin Transformer implemented in MONAI (Cardoso et al., 2022), and our HLIP visual encoder. Since Swin is not compatible with multi-scan inputs, the experiment is conducted on a single-channel 3D input of shape (B,1,224,224,224), where B=1 denotes the batch size. Both ViT and HLIP use a ViT-Base architecture with patch_size=(16,16,16). The Swin Transformer is configured with dim=128, patch_size=(4,4,4), window_size=(7,7,7), depth=(2,2,12,2), and head=(4,8,16,32). Following the original Swin Transformer (Liu et al., 2021) while matching the model size.
@@ E.2 Experiments
model size (M) throughput (img/s; $\uparrow$=better) memory (G; $\downarrow$=better)
@@ Appendix F Discussion
Zero-shot Transferability. We observe that zero-shot transferability does not always correlate with the number of training studies, in contrast to the findings of Udandarao et al. (2024). For example, although keyword search yields more patients with meningioma or metastasis than with glioma, zero-shot performance on glioma is substantially higher. We hypothesize that this discrepancy arises from fundamental differences between natural image and medical imaging datasets. For natural image datasets such as ImageNet, the relationship between the number of positive instances in the training corpus and zero-shot performance is approximately log-linear. This observation implicitly assumes a small or near-zero Bayes error rate, for example, when differentiating boats from dogs. In contrast, in medical imaging, diagnosing meningioma and metastasis is challenging for multifaceted reasons, including lesion size, radiographic features, and anatomical location, which leads to a higher error rate compared to glioma.
@@
Dataset Release Policy. The assets introduced in this work, including a new brain MRI benchmark for zero-shot classification, an effective language–image pre-training implementation for 3D medical imaging, the pre-training recipe, and the model checkpoints, will be fully publicly available. However, we are unable to release the BrainMRI220K and HeadCT240K datasets alongside this work due to privacy constraints associated with neuroimaging data. We plan to release a synthetic dataset in the future, where the generative model is trained on our large-scale dataset.